# Mpemba meets Newton: Exploring the Mpemba and Kovacs effects in the time-delayed cooling law

### v1

@@ V Conclusions
Time delay $\tau$ Free parameter
Maximum time delay $\tau_{\max}$ $e^{-1}\simeq 0.368$
@@
A pivotal role is played by the $\tau$-exp function $\mathcal{E}(t)$, functioning as the time-delay analog of the conventional decaying exponential. Imposing the physical condition that $\mathcal{E}(t)>0$ for all time sets an upper bound, $\tau<\tau_{\max}=e^{-1}$, for the time delay.
As the simplest protocol for the observation of the Mpemba effect, we have assumed that samples A and B were thermalized in the past ($t<-t_{\text{w}}$) to the temperatures ($T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$) of a hot and a cold bath, respectively. Sudden quenches at $t=-t_{\text{w}}$ are then applied to both samples by exchanging their baths, followed by a second quench to temperature $T_{\text{b}}^{\text{c}}$ at $t=0$ for sample B. Notably, the time delay enhances the cooling of sample A while inhibiting that of sample B, leading to the Mpemba effect under specific conditions ($t_{\text{w}}^{\min}<t_{\text{w}}<t_{\text{w}}^{\max}$). The effect is independent of the bath temperatures ($T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$) and is determined by a difference function $\Delta(t)$ parameterized by the control parameters $\tau$ and $t_{\text{w}}$. In the inverse Mpemba effect, the protocol is identical, except for the exchange $T_{\text{b}}^{\text{h}}\leftrightarrow T_{\text{b}}^{\text{c}}$, so that the difference function $\Delta(t)$ is the same as in the direct case. Therefore, the time-delayed Newton’s cooling law with a constant coefficient of heat transfer fails to capture the known fact that heating is faster than cooling [100, 101, 70].
@@
In conclusion, the time-delayed Newton’s law of cooling emerges as a simplistic yet effective phenomenological model that accommodates memory effects. The exploration of the Mpemba and Kovacs effects within its framework has unveiled significant insights into the intricate nature of these memory phenomena. This paper may contribute to a comprehensive analysis of the Mpemba and Kovacs effects, shedding light on their underlying mechanisms and expanding their relevance to diverse systems. The findings presented here not only deepen our understanding of these memory phenomena but also offer valuable insights applicable across various scientific domains, from physics to materials science. The exploration of time-delayed cooling laws paves the way for future research avenues, inviting further investigation into the fascinating interplay between thermal history, memory effects, and complex system behaviors.
@@ Acknowledgements.
The author acknowledges financial support from Grant No. PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033, and from Grant No. IB20079 funded by Junta de Extremadura (Spain) and by ERDF “A way of making Europe.”

### v2

@@ V Conclusions
Delay time $\tau$ Free parameter
Maximum delay time $\tau_{\max}$ $e^{-1}\simeq 0.368$
@@
A pivotal role is played by the $\tau$-exp function $\mathcal{E}(t)$, functioning as the delay-time analog of the conventional decaying exponential. Imposing the physical condition that $\mathcal{E}(t)>0$ for all time sets an upper bound, $\tau<\tau_{\max}=e^{-1}$, for the delay time.
As the simplest protocol for the observation of the Mpemba effect, we have assumed that samples A and B were thermalized in the past ($t<-t_{\text{w}}$) to the temperatures ($T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$) of a hot and a cold bath, respectively. Sudden quenches at $t=-t_{\text{w}}$ are then applied to both samples by exchanging their baths, followed by a second quench to temperature $T_{\text{b}}^{\text{c}}$ at $t=0$ for sample B. Notably, the delay time enhances the cooling of sample A while inhibiting that of sample B, leading to the Mpemba effect under specific conditions ($t_{\text{w}}^{\min}<t_{\text{w}}<t_{\text{w}}^{\max}$). The effect is independent of the bath temperatures ($T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$) and is determined by a difference function $\Delta(t)$ parameterized by the control parameters $\tau$ and $t_{\text{w}}$. In the inverse Mpemba effect, the protocol is identical, except for the exchange $T_{\text{b}}^{\text{h}}\leftrightarrow T_{\text{b}}^{\text{c}}$, so that the difference function $\Delta(t)$ is the same as in the direct case. Therefore, the time-delayed Newton’s cooling law with a constant coefficient of heat transfer fails to capture the known fact that heating is faster than cooling [116, 117, 71].
In general, the Mpemba effect necessitates the consideration of at least three distinct temperatures: the initial temperatures, $T_{\text{A}}(0)$ and $T_{\text{B}}(0)$, of both samples, and the temperature, $T_{\text{b}}^{\text{c}}$, of the shared final bath. The protocol outlined in Fig. 6(a) is relatively straightforward, involving only two bath temperatures ($T_{\text{b}}^{\text{h}}$ and $T_{\text{b}}^{\text{c}}$), along with a waiting time ($t_{\text{w}}$). However, more complex protocols can be devised. For instance, sample B might initially equilibrate to a temperature $T_{\text{b,B}}^{-}$ (with $T_{\text{b}}^{\text{c}}\leq T_{\text{b,B}}^{-}<T_{\text{b}}^{\text{h}}$) for $t<-t_{\text{w}}$, then be quenched at $t=-t_{\text{w}}$ to a bath at temperature $T_{\text{b,B}}^{-*}$ (with $T_{\text{b,B}}^{-}<T_{\text{b,B}}^{-*}\leq T_{\text{b}}^{\text{h}}$), and finally quenched at $t=0$ to the cold bath at temperature $T_{\text{b}}^{\text{c}}$. Moreover, the introduction of two waiting times instead of just one could be contemplated, providing flexibility to accommodate different experimental protocols.
@@
It must be emphasized that. for simplicity, the analysis carried out in this paper has assumed a constant coefficient of heat transfer $\lambda$. However, factors such as dependence on the time-dependent temperature, variation in material properties with temperature, or changing boundary layer conditions due to natural convection might cause $\lambda$ to slightly deviate over time. While a constant $\lambda$ provides a valuable first approximation, acknowledging this potential time-dependence is important for a more comprehensive understanding of the Mpemba and Kovacs effect scenarios. Future investigations could explore the influence of a non-constant $\lambda$ on the results and potentially refine the model for greater accuracy.
In conclusion, the time-delayed Newton’s law of cooling emerges as a simplistic yet effective phenomenological model that accommodates memory effects. The exploration of the Mpemba and Kovacs effects within its framework has unveiled significant insights into the intricate nature of these memory phenomena. This paper may contribute to a comprehensive analysis of the Mpemba and Kovacs effects, shedding light on their underlying mechanisms and expanding their relevance to diverse systems. The findings presented here not only can contribute to deepening our understanding of these memory phenomena, but also offer valuable insights applicable across various scientific domains, from physics to materials science. The exploration of time-delayed cooling laws paves the way for future research avenues, inviting further investigation into the fascinating interplay between thermal history, memory effects, and complex system behaviors.
@@ Acknowledgements.
The author acknowledges financial support from Grant No. PID2020-112936GB-I00 funded by MCIN/AEI/10.13039/501100011033, and from Grant No. IB20079 funded by Junta de Extremadura (Spain) and by European Regional Development Fund (ERDF) “A way of making Europe.”